# Comprehensive Evaluation on Lexical Normalization: Boundary-Aware Approaches for Unsegmented Languages

### v1

@@ Abstract
Lexical normalization research has sought to tackle the challenge of processing informal expressions in user-generated text, yet the absence of comprehensive evaluations leaves it unclear which methods excel across multiple perspectives. Focusing on unsegmented languages, we make three key contributions: (1) creating a large-scale, multi-domain Japanese normalization dataset, (2) developing normalization methods based on state-of-the-art pre-trained models, and (3) conducting experiments across multiple evaluation perspectives. Our experiments show that both encoder-only and decoder-only approaches achieve promising results in both accuracy and efficiency.
@@ 1 Introduction
User-generated text (UGT) is invaluable textual content produced by users’ activities on web platforms such as review sites and social media. UGT’s informality—its frequent use of colloquial expressions—poses substantial challenges to accurate analysis in natural language processing applications. Consequently, researchers have explored lexical normalization (LN), the task of converting non-standard word forms into standard ones. LN has been actively studied particularly for space-delimited languages such as European languages Baldwin et al. (2015); van der Goot et al. (2021). In this study, we investigate LN for unsegmented languages, focusing primarily on Japanese.
@@
Our evaluation reveals three main findings. First, compact encoder-only models deliver the highest inference throughput, and decoder-only models excel in normalization recall—while both yield high normalization precision. Second, training models on 4k–8k sentences yields reasonable precision of around 0.7, and cutting-edge decoder-only models deliver superior recall even with fewer instances. Third, domains rich with unknown informal words exhibit low performance, especially the typo-correction domain. [^1] Our dataset and code will be available at {URL}.
@@ 2 Related Work
To achieve robust word segmentation (WS) for UGT, previous studies have often addressed LN jointly with WS, as exemplified by research on Japanese Sasano et al. (2013); Kaji and Kitsuregawa (2014); Saito et al. (2014, 2017); Higashiyama et al. (2021a), Chinese Wang et al. (2013); Qian et al. (2015), and Thai Haruechaiyasak and Kongthon (2013). Some recent studies have adopted Transformer masked language models (MLMs) Ueda et al. (2023); Pankam et al. (2023).
@@
Regarding evaluation domains, many studies have focused on building short message and social media datasets for European Choudhury et al. (2007); Han and Baldwin (2011); Baldwin et al. (2015); Plank et al. (2020); van der Goot et al. (2021) and Asian languages Kaji and Kitsuregawa (2014); Limkonchotiwat et al. (2021); Nguyen et al. (2024); Kondo et al. (2025), which has led to extensive model development and evaluation in these domains. Some studies have focused on other domains Higashiyama et al. (2021b); Bikaun et al. (2024). However, cross-domain evaluation research covering three or more domains remains scarce.
@@ 3 Japanese Dataset Construction
We followed Higashiyama et al. (2021b)’s annotation criteria; annotation information includes word boundaries based on the short unit word criterion Maekawa et al. (2014) and word attributes such as part-of-speech, lemma, predefined word categories (details in Appendix A.3), and standard forms of non-standard words. [^2] Non-standard forms are those with distinctive orthographic features whose frequency in the reference corpus falls below a threshold, whereas standard forms are those whose frequency exceeds a threshold. See details in Appendix A.4.
@@
As data preparation, the first author extracted sentences with a reasonable length (10–300 characters) from each original 14 datasets and divided the sentences into two candidate sets: a random (Rand) set and a manually curated (Cur) set. Then, four experienced annotators, including an annotator manager, at a data annotation company performed the annotation process as follows.
@@
Sentence selection: From the Rand-set of each original dataset, annotators sequentially selected sentences unless they contained ethically problematic contents or were unclear in meaning. From each Cur-set, annotators intentionally selected sentences containing UGT-specific category words (Appendix A.3).
@@
Word information annotation: Annotators annotated the selected sentences with word information by modifying auto-analyzed results by a morphological analyzer, MeCab Kudo et al. (2004) with UniDic Den (2009).333cwj-3.1.0 (https://clrd.ninjal.ac.jp/unidic/)
@@
Standard form annotation: We adopted the separated steps for standard form (SForm) annotation—assigning an SForm ID for each non-standard word and associating a set of valid standard forms with each SForm ID, where the later step for all sentences was performed by the annotation manager.
@@
During the annotation process, manager A selected a total of 240 sentences from Cur-sets, and then two annotators—B and either C or D—independently annotated those sentences with boundaries and attributes. Inter-annotator agreement (IAA) for non-standard word recognition on these sentences, as measured by F1 score, was 0.836 (See Appendix A.6). [^4] Plank et al. (2020) and van der Goot et al. (2020) used Cohen’s kappa to evaluate IAA in informal word classification, based on whether each given word is normalized or not. This metric is not directly applicable to unsegmented languages. The datasets’ high annotation consistency and large size suggest its usefulness, and this is further demonstrated by the experiments in §6, where the evaluated models achieve high normalization accuracy.
@@ 4 Task Definition
Following previous studies Sasano et al. (2013); Baldwin et al. (2015), we define LN as a task of boundary-aware span extraction and conversion, in which a system not only generates a normalized text but also identifies the original spans of each informal words (or phrases). This task can be performed on any LN dataset, provided that informal-to-formal alignments are annotated. [^5] Otherwise, a non-trivial alignment processing is needed. We provide an alignment examples in Appendix A.5. Unlike the text-to-text conversion task Ikeda et al. (2016), the boundary-aware task enables fine-grained evaluation at the normalization span level and offers better interpretability of system outputs. An example input-output pair for our task is shown in Figure 1.
Formally, an LN system takes as input a source sentence, namely, a sequence of $n$ character (or subword) tokens $\bm{x}=\bm{x}_{0:n}=[x_{0},\dots,x_{n-1}]$, and is required to predict the set of non-standard word spans and their standard forms $S=\{(b,e,s)\}$. Here, $(b,e)$ ($0\!\leq\!b\!\leq\!e\!\leq\!n$) indicates a span of an non-standard word $x_{b:e}$ with length $e-b$ in the source sentence, and $s$ indicates its standard form. Each standard form $s$ is a string with length $\geq 0$,
where length $=0$ indicates that the non-standard word should be deleted in the normalized sentence. When $b=e$, a zero-length span indicates some token(s) should be inserted into the position $b$.
Figure 1: Flow of our detect&infill approach for an input text “{CJK}UTF8ipxmついったみてる,” which means “(I’m) looking at Twitter.” “M” and “S” represent the MASK and SEP token, respectively. The original characters “{CJK}UTF8ipxmついった” follow the SEP token, but are omitted in the Figure.
@@ 5.1 Infilling Approach
We refer to the above method as the Full-Seg approach. [^7] We report preliminary experiments on additional variants in the Appendix B During training, the model is optimized using the sum of cross-entropy losses over multiple subtasks from both steps.
@@ 5.2 Generative Approaches
Plain {CJK}UTF8ipxmツイッターみてる
Struct {CJK}UTF8ipxm[[ついった> >ツイッター]]みてる
Span {CJK}UTF8ipxmついった> >ツイッター> > 0
Table 1: Expected output text of each generative approach for an input text “{CJK}UTF8ipxmついったみてる.”
@@
Lang Dataset Set #Sent #Norm
ja JMLN train 13,196 5,879
dev 1,880 791
test-C 3,786 1,705
test-R 2,540 510
th VISTEC-2021 train 40,000 130,790
test 10,000 32,819
@@ 6 Experiments
To address these EQs, we evaluated various LN methods using our Japanese dataset and an existing Thai dataset. Specifically, experiments in §7.17.3, §7.4, and §7.57.6 correspond to EQ1, EQ2, and EQ3, respectively.
As evaluation metrics, we use precision and recall at the normalization-span level for individual non-standard words, as well as the F0.5 score, which emphasizes precision over recall (see details in Appendix C.1).
@@
Table 3: JMLN test results of Japanese LN models (E: encoder, S: seq2seq, D: decoder). $\hookrightarrow$” indicates the continual pre-trained model derived from the base model listed in the previous row. The best score within each size group is shown in bold. For each backbone model, the better of the Struct and Span approaches is underlined. Scores where the Span approach shows a +5%, +10%, or -10% increase/decrease compared to the Struct approach are highlighted with $\blacksquare$ blue, $\blacksquare$ light blue, and $\blacksquare$ pink backgrounds, respectively.
@@ 6.1 Datasets
Table 2 shows the statistics of two experimental datasets: JMNL and Thai VISTEC-2021 Limkonchotiwat et al. (2021). [^8] https://github.com/mrpeerat/OSKut For JMLN, we divided the Cur-set for each domain (01–14) into train, dev, and test-C sets, and merged the all train, dev, and test-C sets into unified train, dev, and test-C sets across all domains, respectively. [^9] We created the unified Rand-set across all registers similarly, but omit experimental results on this set. For VISTEC-2021, we followed the provided training/test split and regarded randomly-sampled 5% sentences in the training set as a dev set.
@@ 6.2 Models
As the backbone of LN systems, we used the following Japanese or multilingual pre-trained models in Japanese experiments: BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and DeBERTa He et al. (2021) as character-level encoder-only models, T5 Raffel et al. (2020) and mT5 Xue et al. (2021) as encoder-decoder models, and Llama 3.1/3.2 Grattafiori et al. (2024), Qwen2.5 Qwen Team (2025), Lllama-3.1-Swallow Fujii et al. (2024), TinySwallow Shing et al. (2025) and Sarashina2/2.2 Intuitions (2024a, b) as decoder-only models. Similarly, we used the following pre-trained models in the Thai experiments: RoBERTa Yasuoka (2023) as a character-level encoder-only model, T5 and mT5 as encoder-decoder models, and Llama 3.1/3.2, Qwen2.5, Typhoon 2 Pipatanakul et al. (2024), OpenThaiGPT 1.5 Yuenyong et al. (2025), and SeaLLMs 3 Zhang et al. (2024a) as decoder-only models. Specific model instances are listed in Appendix C.2.
We fine-tuned each model, with applying LoRA Hu et al. (2022) or QLoRA Dettmers et al. (2023) to some decoder-only models, and selected the model checkpoint with the best F0.5 score on the dev set. We fine-tuned all models twice and report mean scores for two runs. The hyperparameter settings are listed in Appendix C.3.
@@ 7.1 Normalization Accuracy for Japanese
We evaluated LN methods with three type of architectures on the JMLN test-C set. Table 3 shows the performance of encoder-only models, and encoder-decoder and decoder-only models with both Struct and Span approaches. [^10] The Plain approach yielded performance similar to the other two approaches on the dev set, as shown in §D.3.
The observed results are as follows. (1) Among encoder-only models, the large models outperformed base models, while different backbone models showed similar performance. (2) Among generative methods, the Span approach basically achieved performance comparable to or better than the Struct approach in many cases. [^11] The only exception is the Qwen2.5 model series; the models generated many nonsensical text flagments—such as “str1> >str2> >0,” where “str1” did not appear in the original input text—until reaching the maximum output length, resulted in low precision and long inference time. (3) Within the same model series—T5, mT5, Llama-3.2, Qwen2.5, and Sarashina2.2—performance improved with increasing model size up to 8B (See Figure in Appendix D.1). However, in our preliminary experiment on the dev set, we observed no salient additional gains from the 13B–15B models. (4) Performance within groups of similarly sized models was not equivalent; certain series—specifically Sarashina2.2 series, followed by Swallow, demonstrated saliently superior performance. (5) These strong decoder-only models outperformed models with the other two architectures in recall.
@@ 7.2 Normalization Accuracy for Thai
7.3 Inference Throughput
For selected models with high normalization accuracy, we measured their inference throughput using the JMLN test set, on both an NVIDIA V100 GPU with 32 GiB memory and an H200 GPU with 140 GiB memory (See detailed settings in Appendix C.6). Table 5 shows the results.
@@ 7.3 Inference Throughput
09  NC-VID Video desc. 0.39 0.645 0.627 0.538 0.783 0.677 0.730 0.629
10  NC-PED Forum 0.27 0.795 0.850 0.747 0.666 0.492 0.808 0.772

### v2

@@ Abstract
Lexical normalization research has sought to tackle the challenge of processing informal expressions in user-generated text, yet the absence of comprehensive evaluations leaves it unclear which methods excel across multiple perspectives. Focusing on unsegmented languages, we make three key contributions: (1) creating a large-scale, multi-domain Japanese normalization dataset, (2) developing normalization methods based on state-of-the-art pre-trained models, and (3) conducting experiments across multiple evaluation perspectives. Our experiments show that both encoder-only and decoder-only approaches achieve promising results in both accuracy and efficiency. [^1] Our dataset and code will be available at https://github.com/shigashiyama/jmln.
@@ 1 Introduction
User-generated text (UGT) is invaluable textual content produced by users’ activities on web platforms such as review sites and social media. UGT’s informality—its frequent use of colloquial expressions—poses substantial challenges to accurate analysis in natural language processing applications. Consequently, researchers have explored lexical normalization (LN), the task of converting non-standard word forms into standard ones. LN has been actively studied particularly for space-delimited languages such as European languages Baldwin et al. (2015); van der Goot et al. (2021). In this study, we investigate LN for unsegmented languages in writing, focusing primarily on Japanese.
@@
Our evaluation reveals three main findings. First, compact encoder-only models deliver the highest inference throughput, and decoder-only models excel in normalization recall—while both yield high normalization precision. Second, training models on 4k–8k sentences yields reasonable precision of around 0.7, and cutting-edge decoder-only models deliver superior recall even with fewer instances. Third, domains rich with unknown informal words exhibit low performance, especially the typo-correction domain.
@@ 2 Related Work
To achieve robust word segmentation (WS) for UGT, previous studies have often addressed LN jointly with WS, as exemplified by research on Japanese Sasano et al. (2013); Kaji and Kitsuregawa (2014); Saito et al. (2014, 2017); Higashiyama et al. (2021a), Chinese Wang et al. (2013); Qian et al. (2015), and Thai Haruechaiyasak and Kongthon (2013). Some recent studies have adopted Transformer masked language models (MLMs) with a text-editing approach Ueda et al. (2023) and a two-step approach—first detecting informal tokens and then predicting formal tokens Pankam et al. (2023).
@@
Regarding evaluation domains, many studies have focused on building short message and social media datasets for European Choudhury et al. (2007); Han and Baldwin (2011); Baldwin et al. (2015); Plank et al. (2020); van der Goot et al. (2021) and Asian languages Kaji and Kitsuregawa (2014); Limkonchotiwat et al. (2021); Nguyen et al. (2024); Kondo et al. (2025), which has led to extensive model development and evaluation in these domains. Some studies have focused on other domains such as blog and Q&A site Higashiyama et al. (2021b), and maintenance short text Bikaun et al. (2024). However, cross-domain evaluation research covering three or more domains remains scarce.
@@ 3 Japanese Dataset Construction
We followed Higashiyama et al. (2021b)’s annotation criteria. The annotation information includes word boundaries based on the short-unit word criterion Maekawa et al. (2014), as well as word attributes such as part-of-speech, lemma, predefined word categories (details in Appendix A.4), and standard forms of non-standard words. [^2] Non-standard forms are those with distinctive orthographic features whose frequency in the reference corpus falls below a threshold, whereas standard forms are those whose frequency exceeds a threshold. See details in Appendix A.5. Thus, the dataset supports both LN and Japanese morphological analysis, which involves word boundary detection, part-of-speech tagging, and lemmatization. Nevertheless, this study focuses on the evaluation of LN, leaving morphological analysis tasks outside its scope.
@@
As part of data preparation, we extracted sentences with a reasonable length (10–300 characters) as a candidate set from each of the original 14 datasets. Four experienced annotators, one of whom was a manager, at a data annotation company then carried out the annotation process as follows.
@@
Sentence selection: From each candidate set for each original dataset, annotators intentionally selected sentences containing UGT-specific category words (Appendix A.4).
@@
Word information annotation: Annotators annotated the selected sentences with word information by modifying auto-analyzed results by a morphological analyzer, MeCab Kudo et al. (2004) with UniDic (unidic-cwj-3.1.0Den (2009). [^3] https://clrd.ninjal.ac.jp/unidic/
@@
Standard form annotation: We adopted separated steps for standard form (SForm) annotation—(1) recognizing non-standard words by assigning variant-form type categories (Appendix A.4), (2) assigning an SForm ID to each non-standard word, and (3) associating a set of valid standard forms with each SForm ID. While the first step was performed by annotators, the second and third steps were performed by the annotation manager.
@@
During the annotation process, manager A selected a total of 240 sentences from the candidate sets, and then two annotators—B and either C or D—independently annotated those sentences with boundaries and attributes. Inter-annotator agreement (IAA) for non-standard word recognition on these sentences, as measured by F1 score, was 0.836 (see Appendix A.7). [^4] Plank et al. (2020) and van der Goot et al. (2020) used Cohen’s kappa to evaluate IAA in informal word classification, based on whether each given word is normalized or not. This metric is not directly applicable to unsegmented languages. The datasets’ high annotation consistency and large size suggest its usefulness, and this is further demonstrated by the experiments in §6, where the evaluated models achieve high normalization accuracy.
@@ 4 Task Definition
Following previous studies Sasano et al. (2013); Baldwin et al. (2015), we define LN as a task of boundary-aware span extraction and conversion, in which a system not only generates a normalized text but also identifies the original spans of each informal words (or phrases). An example input-output pair for our task is shown in Figure 1.
This task can be carried out independently of the tokenization unit and applied to any LN dataset, as long as informal-to-formal alignments are annotated. [^5] Otherwise, a non-trivial alignment processing is needed. We provide an alignment examples in Appendix A.6. In contrast to the text-to-text conversion task Ikeda et al. (2016), the boundary-aware task allows fine-grained evaluation at the normalization-span level and provides better interpretability of system outputs. This is because span-level evaluation metrics directly assess the validity of each normalization instance, leading to more robust evaluation across sentences with varying densities of non-standard tokens. Further discussion on the task definition is provided in Appendix B
Formally, an LN system takes as input a source sentence, namely, a sequence of $n$ character (or subword) tokens $\bm{x}=\bm{x}_{0:n}=[x_{0},\dots,x_{n-1}]$, and is required to predict the set of non-standard word spans and their standard forms $P=\{(b,e,s)\}$. Here, $(b,e)$ ($0\!\leq\!b\!\leq\!e\!\leq\!n$) indicates a span of an non-standard word $x_{b:e}$ with length ebe-b in the source sentence, and $s$ indicates its standard form. [^6] A gold-standard normalization instance $i$ can have multiple standard forms and is thus represented as $(b_{i},e_{i},\{s_{i,k}\}_{k})$. Each standard form $s$ is a string with length $\geq 0$,
where length =0=0 indicates that the non-standard word should be deleted in the normalized sentence. When $b=e$, a zero-length span indicates some token(s) should be inserted into the position $b$.
@@ 5.1 Infilling Approach
Figure 1: Flow of our detect&infill approach for an input text “ついったみてる,” which means “(I’m) looking at Twitter.” “M” and “S” represent the MASK and SEP token, respectively. The original characters “ついった” follow the SEP token, but are omitted in the Figure.
@@
We refer to the above method as the Full-Seg approach. During training, the model is optimized using the sum of cross-entropy losses over multiple subtasks from both steps.
As a variant of this approach, the detection step can be performed eigher with full-word segmentation or, alternatively, with partial word segmentation, which requires only the boundaries of (continuous) non-standard words in a sentence. The latter approach, referred to as Part-Seg, is introduced in Appendix C. We report a comparison among variants, including these two, in Appendix E.3.
@@ 5.2 Generative Approaches
Plain ツイッターみてる
Struct [[ついった> >ツイッター]]みてる
Span ついった> >ツイッター> > 0
Table 1: Expected output text of each generative approach for an input text “ついったみてる.”
@@ 6 Experiments
To address these EQs, we evaluated various LN methods using our Japanese dataset and an existing Thai dataset. Specifically, experiments in §7.17.3 (and §7.6) correspond to EQ1, §7.4 to EQ2, and §7.5 and §7.7 to EQ3.
We use precision, recall, and F0.5 score at the normalization-span level for individual non-standard words as evaluation metrics. While normalized results with low recall do not harm downstream task performance compared to original texts, low-precision outputs introduce spurious tokens that can degrade downstream accuracy. We therefore adopt the F0.5 score, which emphasizes precision over recall, similarly to other text correction tasks Wang et al. (2021). See Appendix D.1 for details of the score calculations.
Lang Dataset Set #Sent #Norm
ja JMLN train 13,196 5,879
dev 1,880 791
testc 3,786 1,705
testr 2,540 510
th VISTEC-2021 train 40,000 130,790
test 10,000 32,819
@@
Table 3: JMLN test results of Japanese LN models (E: encoder, S: seq2seq, D: decoder). $\hookrightarrow$” indicates the continual pre-trained model derived from the base model listed in the previous row. For each size group (separated by horizontal lines), the best score is indicated in bold. For each backbone model, the better of the Struct and Span approaches is underlined. Scores where the Span approach shows a +5%, +10%, or -10% increase/decrease compared to the Struct approach are highlighted with $\blacksquare$ blue, $\blacksquare$ light blue, and $\blacksquare$ pink backgrounds, respectively.
@@ 6.1 Datasets
Table 2 shows the statistics of two experimental datasets: JMLN and Thai VISTEC-2021 Limkonchotiwat et al. (2021). [^8] https://github.com/mrpeerat/OSKut/tree/main/VISTEC-TP-TH-2021 For JMLN, we divided the annotation data for each domain (01–14) into train, dev, and testc sets, and merged the all train, dev, and testc sets into unified train, dev, and testc sets across all domains, respectively. [^9] We also created an additional test set, the testr set, but focus on experiments with the testc set. See Appendix A.1 for more details on the testr set. Unless otherwise stated, we simply refer to the testc set as the “test set.” For VISTEC-2021, we followed the provided training/test split and regarded randomly-sampled 5% sentences in the training set as a dev set.
@@ 6.2 Models
As the backbone of LN systems, we used the following Japanese or multilingual pre-trained models in Japanese experiments: BERT Devlin et al. (2019), RoBERTa Zhuang et al. (2021), and DeBERTa He et al. (2021) as character-level encoder-only models, T5 Raffel et al. (2020) and mT5 Xue et al. (2021) as encoder-decoder models, and Llama 3.1/3.2 Grattafiori et al. (2024), Qwen2.5 Qwen Team (2025), Lllama-3.1-Swallow Fujii et al. (2024), TinySwallow Shing et al. (2025) and Sarashina2/2.2 Intuitions (2024, 2025) as decoder-only models. [^10] Languages supported by each model instance are shown in Table 23. Notably, the encoder-only models used in this study tokenize at the character level, whereas the other models employ subword-level tokenization.
Similarly, we used the following Thai or multilingual pre-trained models in the Thai experiments: RoBERTa Yasuoka (2023) as a character-level encoder-only model, T5 and mT5 as encoder-decoder models, and Llama 3.1/3.2, Qwen2.5, Typhoon 2 Pipatanakul et al. (2024), OpenThaiGPT 1.5 Yuenyong et al. (2025), and SeaLLMs 3 Zhang et al. (2024a) as decoder-only models. Specific model instances are listed in Appendix D.2.
We fine-tuned each model, with applying LoRA Hu et al. (2022) or QLoRA Dettmers et al. (2023) to some decoder-only models, and selected the model checkpoint with the best F0.5 score on the dev set. We fine-tuned all models twice with different random seeds and report mean scores for two runs. The hyperparameter settings are listed in Appendix D.3.
@@ 7.1 Normalization Accuracy for Japanese
We evaluated LN methods with three types of architectures on the JMLN test set. Table 3 shows the performance of encoder-only models, and encoder-decoder and decoder-only models with both Struct and Span approaches. [^11] The Plain approach yielded performance similar to the other two approaches on the dev set, as shown in §E.4.
The observed results are as follows. (1) Among encoder-only models, the large models outperformed base models, while different backbone models showed similar performance. (2) Among generative methods, the Span approach basically achieved performance comparable to or better than the Struct approach in many cases. [^12] The only exception is the Qwen2.5 model series; the models generated many nonsensical text flagments—such as “str1> >str2> >0,” where “str1” did not appear in the original input text—until reaching the maximum output length, resulted in low precision and long inference time. (3) Within the same model series—T5, mT5, Llama-3.2, Qwen2.5, and Sarashina2.2—performance improved with increasing model size up to 8B (see Figure in Appendix E.1). However, in our preliminary experiment on the dev set, we observed no salient additional gains from the 13B–15B models. (4) Performance within groups of similarly sized models was not equivalent; certain series—specifically Sarashina2.2 series, followed by Swallow, demonstrated saliently superior performance. (5) These strong decoder-only models outperformed models with the other two architectures in recall.
@@ 7.2 Normalization Accuracy for Thai
7.3 Inference Throughput
For selected models with high normalization accuracy, we measured their inference throughput using the JMLN test set, on both an NVIDIA V100 GPU with 32 GiB memory and an H200 GPU with 140 GiB memory (see detailed settings in Appendix D.6). Table 5 shows the results.
@@ 7.3 Inference Throughput
Figure 2: JMLN test results for each training data size.
@@
09  NC-VID Video desc. 0.39 0.645 0.627 0.538 0.666 0.492 0.730 0.629
10  NC-PED Forum 0.27 0.795 0.850 0.747 0.783 0.677 0.808 0.772
@@
Backbone Approach Det Norm Det-Norm
P R F0.5 P R F0.5 P R
DeBERTa-large Full-Seg 0.865 0.655 0.813 0.750 0.568 0.705 0.115 0.087
T5-large Span 0.801 0.597 0.750 0.704 0.525 0.659 0.097 0.072
Sarashina2.2-3B Span 0.829 0.700 0.800 0.781 0.660 0.754 0.048 0.041
Table 7: JMLN test results of representative three models for detection (Det) and normalization (Norm) tasks. The best score among three models is highlighted in bold.